How many grams of active ingredients are contained in the following: 50 ml of a 75% solution?

3 fl. oz. of a 12% solution?

2 teaspoonful of a 27.5% solution?

0.5 ml of a 20% solution?
Answer #1

A. 75% means 75gm in 100 ml

then in 50 ml it will be 75/2= 37.5 gm

B. 3 fl.oz.=88.72 ml

12% means 12 gm in 100 ml

means 100 ml contains= 12 gm

1ml will= 12/100

88.72 ml will have =(12/100) X 88.72

=10.645 gm

C. 2 teaspoonful of a 27.5% solution?

2 teaspoons=9.857 ml

27.5% means 27.5 gm in 100 ml

in 9.875 ml it will be= (27.5/100) X 9.875

=2.72 gm

D. 0.5 ml of a 20% solution?

20 % means 20 gm in 100 ml

in 1ml = 20/100=0.2 gm

hence in 0.5 ml it will be= 0.2/2

=0.1 gm